### v1

@@ 5.1 Experimental Settings
All experiments were conducted on a single NVIDIA A100 GPU. We employed random search to tune the hyperparameters of WAVE++, while the baseline methods used the configurations reported by Zhao et al. (2022), allowing for fair comparisons. Specifically, we froze the parameters of the pre-trained BERT model and optimized only the MLP-based relation classifiers and the prompt pools. Under these conditions, training WAVE++ took approximately seven hours on the FewRel dataset and one hour on the TACRED dataset.
@@ 5.2 Main Results
Figure 5: Variation in average accuracy (%) for individual tasks during the training process using WAVE++, WAVE-CRE, and EoE on the FewRel dataset.
In Table 2, we present a detailed performance comparison of WAVE++ against both rehearsal-free and rehearsal-based CRE methods. We first benchmarked WAVE++ against other rehearsal-free approaches, finding that it consistently outperforms all baselines. Notably, on the FewRel dataset at the final task, WAVE++ exceeds the performance of the second-best method, EoE, by 2%. Furthermore, WAVE++ surpasses all other methods across various stages of training, as demonstrated in Figure 5, where the accuracy of WAVE++ progressively outperforms EoE during training. Although WAVE++ shows minor dips relative to EoE at certain points on TACRED, its performance on the final tasks, as well as in scenarios with an increasing number of tasks, ultimately reflects its superior resistance to catastrophic forgetting.
In addition, WAVE++ demonstrates substantial improvements over its predecessor, WAVE-CRE, throughout the training process. By the final task, WAVE++ achieves an accuracy that surpasses WAVE-CRE by approximately 2% on FewRel and 4% on TACRED. The accuracy across individual tasks also shows significant improvement over the course of training, as depicted in Figure 5. These gains can largely be attributed to the incorporation of label descriptions and an enhanced voting-based task prediction mechanism, as opposed to the previous MLP-based approach. Further details on these improvements will be discussed in Section 5.3.
Furthermore, we evaluate WAVE++ against recent state-of-the-art, rehearsal-based CRE methods. Despite not utilizing a memory buffer for replay, WAVE++ outperforms all evaluated rehearsal-based methods. On both datasets, WAVE++ demonstrates a significant performance advantage and exhibits robust resistance to catastrophic forgetting, even in the absence of prior sample storage. These findings underscore the effectiveness and efficiency of WAVE++ as a robust solution for continual relation extraction.
@@ 5.3 Ablation Studies
Task-Specific Prompt Pool
To assess the impact of the task-specific prompt pool on the performance of WAVE++, we conducted experiments under two conditions: WAVE++ with a prompt pool and WAVE++ with only a single task-specific prompt per task. The detailed results are presented in Table 3. The inclusion of a prompt pool enhances performance across both datasets. As discussed in Section 4.1, the prompt pool enables the dynamic selection of prefix experts within each task, rather than relying on a fixed set. This flexibility allows the model to better capture task-specific variations, leading to improved performance.
@@
To comprehensively evaluate the impact of label descriptions, we conducted experiments comparing EoE and our proposed WAVE++ model under two conditions: one with label descriptions and one without. As presented in Table 4, the removal of label descriptions results in a reduction of approximately 2% in the average accuracy of WAVE++ across both datasets. The inclusion of label descriptions provides additional contextual information, thereby facilitating more effective and robust learning of relational features, which in turn enhances classification performance. These findings emphasize the important role of label descriptions in capturing essential relational characteristics and improving within-task predictions.
Additionally, label descriptions significantly improve the accuracy of EoE on the TACRED dataset, as shown in Table 5. This further demonstrates the effectiveness of label descriptions as a valuable component in enhancing continual relation extraction performance.
@@
Number of Label Descriptions $D$
To explore the impact of the number of label descriptions $D$ in Equation (18) on the performance of WAVE++, we conducted experiments using the FewRel and TACRED datasets with varying values of $D$. Specifically, we tested configurations with $D$ ranging from zero to seven. The results, presented in Table 6, demonstrate that WAVE++ maintains consistent performance even with a limited number of label descriptions. These findings indicate that WAVE++ is not heavily reliant on a large number of label descriptions for effective performance. In fact, the model achieves high accuracy with only a few descriptions, thus offering a favorable balance between predictive capability and computational efficiency. This characteristic highlights the versatility of WAVE++ for practical scenarios, where acquiring extensive label descriptions may be challenging or costly.
TACRED
$L$=8, $K$=1 97.3 93.3 90.2 87.9 86.0 84.9 84.1 83.2 82.7 82.3
$L$=4, $K$=2 97.3 93.4 90.3 88.0 86.2 85.1 84.2 83.3 82.8 82.2
$L$=2, $K$=4 97.5 93.4 90.5 88.1 86.2 85.2 84.0 83.5 82.7 82.1
$L$=1, $K$=8 97.6 93.6 90.7 88.2 86.4 85.4 84.3 83.7 83.2 82.5
Number of Experts per Prompt $L$
Similarly, we investigate the impact of varying the prompt length $L$ on model performance. The number of prompts selected, $K$, was adjusted to maintain a consistent total number of experts across experiments, ensuring a fair comparison. The results are summarized in Table 7. As discussed in Section 4.1, setting $L=1$ provides greater flexibility in expert selection, thereby enhancing the model’s expressiveness, since each prefix expert is assigned a distinct key within the prompt pool. Empirical results demonstrate that setting $L=1$ yields the best performance compared to other values of $L$.

### v2

@@ 5.1 Experimental Settings
All experiments were conducted on a single NVIDIA A100 GPU. We tuned the hyperparameters for WAVE++ using random search, while for baseline methods, we adopted the configurations reported in Zhao et al. (2022) to ensure a fair comparison. Specifically, we froze the parameters of the pre-trained BERT model and optimized only the MLP-based relation classifiers and the prompt pools. WAVE++ has 3.5 million trainable parameters, approximately 300,000 fewer than the original WAVE model (3.8 million). In practice, the model trains for approximately seven hours on FewRel and one hour on TACRED. Following Zhou et al. (2024), the maximum number of experts allowed in cascade voting is set to $m=2$. A detailed analysis of training and inference time is provided in H.
@@ 5.2 Main Results
In Table 2, we present a detailed performance comparison of WAVE++ against both rehearsal-free and rehearsal-based CRE methods. We first benchmark WAVE++ against other rehearsal-free approaches and find that it consistently outperforms all baselines. Notably, on the final task of the FewRel dataset, WAVE++ exceeds the performance of the second-best method, EoE, by 2.0%. This superiority is consistent across various training stages, as demonstrated in Figure 5. Although WAVE++ exhibits minor performance dips relative to EoE at certain points on TACRED, its strong performance on the final task highlights its superior resistance to catastrophic forgetting.
Figure 5: Average accuracy (%) on individual tasks throughout the continual learning process for WAVE++, WAVE-CRE, and EoE on the FewRel dataset.
WAVE++ also demonstrates substantial improvements over its predecessor, WAVE-CRE, throughout the training process. By the final task, WAVE++ achieves an accuracy that surpasses WAVE-CRE by approximately 2.0% on FewRel and 4.0% on TACRED. The accuracy across individual tasks also shows significant improvement over the course of training, as depicted in Figure 5. These gains can largely be attributed to the incorporation of label descriptions and an enhanced voting-based task prediction mechanism, which replaces the previous MLP-based approach. A detailed ablation study of these components is presented in Section 5.3.
Furthermore, we evaluate WAVE++ against recent state-of-the-art, rehearsal-based CRE methods. Despite not utilizing a memory buffer for replay, WAVE++ outperforms all evaluated rehearsal-based methods. On both datasets, WAVE++ demonstrates a significant performance advantage and exhibits robust resistance to catastrophic forgetting, all without storing prior samples. These findings underscore the effectiveness and efficiency of WAVE++ as a powerful solution for continual relation extraction.
@@ 5.3 Ablation Studies
Effect of Task-Specific Prompt Pool
To assess the impact of the task-specific prompt pool, we compare the full WAVE++ model against an ablated variant restricted to a single prompt per task. The results, presented in Table 3, show that using a prompt pool consistently enhances performance across both datasets. As discussed in Section 4.1, the prompt pool enables the dynamic selection of prefix experts within each task, rather than relying on a single, static set. This flexibility allows the model to better capture task-specific variations, leading to improved performance.
@@
To comprehensively evaluate the impact of label descriptions, we evaluate both our proposed WAVE++ model and the EoE baseline with and without this component. As shown in Table 4, removing label descriptions reduces the average accuracy of WAVE++ by approximately 2.0% across both datasets.This performance gain can be attributed to the additional contextual information provided by the descriptions, which enables the model to learn more robust relational features and, consequently, enhances classification accuracy. These findings underscore the crucial role of label descriptions in improving within-task prediction performance.
Furthermore, we observe that label descriptions also significantly improve the performance of the EoE baseline on the TACRED dataset, as detailed in Table 5. This result confirms that integrating label descriptions is a broadly effective strategy for enhancing performance in continual relation extraction. A further ablation on the impact of the number of descriptions is provided in D.
Effect of Number of Experts per Prompt $L$
We also evaluated the impact of the prompt length $L$ on model performance. To ensure a fair comparison, the number of selected prompts $K$, was adjusted to keep the total number of experts constant across all experiments. The results, summarized in Table 6. As discussed in Section 4.1, setting $L=1$ maximizes selection flexibility by assigning each prefix expert a unique key, thereby enhancing the model’s expressive power. Empirically, this configuration yields the best performance compared to all other tested values of $L$.
Effectiveness of Cascade Voting for Task Identity Inference
We replace the conventional MLP head for task classification with a cascade voting mechanism for task identity prediction. This approach both reduces training overhead and improves task identification by aggregating predictions from all prompt pools. To evaluate its effectiveness, we compare it against the MLP-based method used in WAVE-CRE (Le et al., 2025a). The task classification accuracies are presented in Table 7.
TACRED
LL=8, KK=1 97.3 93.3 90.2 87.9 86.0 84.9 84.1 83.2 82.7 82.3
LL=4, KK=2 97.3 93.4 90.3 88.0 86.2 85.1 84.2 83.3 82.8 82.2
LL=2, KK=4 97.5 93.4 90.5 88.1 86.2 85.2 84.0 83.5 82.7 82.1
LL=1, KK=8 97.6 93.6 90.7 88.2 86.4 85.4 84.3 83.7 83.2 82.5
FewRel
EoE 100 95.3 95.7 94.4 92.5 91.6 88.3 87.5 86.8 86.0
WAVE-CRE 100 96.8 94.5 93.3 91.0 89.5 88.6 87.6 86.4 85.4
WAVE++ 100 97.9 96.1 95.1 94.0 92.7 91.7 90.8 89.6 88.3
TACRED
EoE 100 95.2 87.5 85.0 82.0 82.4 83.3 82.7 84.1 82.4
WAVE-CRE 100 96.8 91.1 86.8 84.9 83.6 82.1 80.4 80.2 79.2
WAVE++ 100 95.9 92.0 90.5 88.7 87.9 86.9 85.9 85.2 84.8
FewRel
WAVE++ w/o GP 98.2 92.2 88.5 84.2 80.2 76.2 72.9 70.2 67.3 62.1
WAVE++ 98.2 95.8 95.1 94.1 92.7 91.9 90.2 89.9 89.0 87.7
TACRED
WAVE++ w/o GP 97.6 93.6 80.3 72.2 70.1 67.2 65.1 63.6 62.1 60.3
WAVE++ 97.6 93.6 90.7 88.2 86.4 85.4 84.3 83.7 83.2 82.5
Table 8: Average accuracy (%) of WAVE++ with and without generative replay. The best result for each dataset and metric is bolded
The results demonstrate that WAVE++ significantly outperforms WAVE-CRE, achieving an accuracy improvement of approximately 3.0% on FewRel and 4.0% on TACRED after the final task. Beyond these accuracy gains, the cascade voting mechanism exhibits superior stability. The MLP-based approach, in contrast, shows occasional sharp declines in accuracy, indicating a sensitivity to distribution shifts when new tasks are introduced. The voting mechanism, however, maintains consistent performance, highlighting its robustness to heterogeneous task distributions. Moreover, WAVE++ also surpasses EoE in task prediction accuracy. These findings underscore the effectiveness and resilience of the cascade voting mechanism in complex, multi-task scenarios where robustness to distribution shifts is crucial.
Effect of Generative Replay
To assess the impact of generative replay, we remove it from the WAVE++ MLP relation classifier (Table 8). The exclusion of generative replay leads to a substantial performance decline across all tasks, a clear indicator of catastrophic forgetting. By the final training stage, overall performance falls by more than 22%. On the FewRel dataset, for instance, accuracy on previously learned tasks drops by up to 4% after the introduction of each new task. Similarly, a sharp decline is observed on the TACRED dataset, where performance drops by 13% between the second and third tasks. These results underscore the critical role of generative replay in mitigating forgetting and preserving knowledge.
6 Conclusion
In this paper, we introduce a novel perspective on prompt-based continual learning by interpreting existing methods through the lens of mixture of experts models. We demonstrated that contemporary prompt-based continual learning approaches can be viewed as a special case of MoE architectures. Building on this insight, we present WAVE++, which shares key conceptual similarities with sparse MoE designs. Our framework addresses several challenges in prompt-based continual learning, including suboptimal task identification, sensitivity to distribution shifts, and difficulties in handling within-task diversity. Specifically, we mitigate these issues by employing a cascade voting mechanism for task prediction and using task-specific prompt pools, which are complemented by label descriptions to enhance continual relation extraction performance. Despite these advancements, hurdles remain in improving the efficiency and effectiveness of WAVE++. Notably, the method heavily depends on the design of prompt pools and experts, suggesting opportunities to refine and optimize these components. While our approach attenuates catastrophic forgetting, retaining knowledge of previous tasks remains challenging, echoing issues highlighted in prior studies. In particular, if prompt pools are not properly applied during testing, forgetting can reemerge. Moreover, although the prompt pools enhance the expressiveness of our method, the current prefix-tuning experts are relatively simplistic. Future work may explore more sophisticated expert architectures to further enhance the model’s capabilities.
Appendix A Proof of Propositions
In this appendix, we provide proofs for key propositions in the main text.
A.1 Proof of Proposition 3.1
Specifically, let ${\bm{X}}=\left[{\bm{x}}_{1}^{\top},\dots,{\bm{x}}_{N}^{\top}\right]^{\top}\in\mathbb{R}^{Nd}$ denote the concatenation of all input token embeddings. We define $E_{j}\in\mathbb{R}^{d\times Nd}$ as a selector matrix such that $E_{j}{\bm{X}}={\bm{x}}_{j}$, thereby extracting the $j$-th token from the input sequence. From Equation (2), the output of the $l$-th attention head can be expressed as follows:
${\bm{h}}_{l}$ $=\mathrm{softmax}\left(\frac{{\mathbf{X}}^{Q}W_{l}^{Q}{W_{l}^{K}}^{\top}{{\mathbf{X}}^{K}}^{\top}}{\sqrt{d_{v}}}\right){\mathbf{X}}^{V}W_{l}^{V}$
$=\left[{\bm{h}}_{l,1},\dots,{\bm{h}}_{l,N}\right]^{\top}\in\mathbb{R}^{N\times d_{v}},$ (18)
${\bm{h}}_{l,i}$ $=\sum_{j=1}^{N}\frac{\exp\left(\frac{{\bm{x}}_{i}^{\top}W_{l}^{Q}{W_{l}^{K}}^{\top}{\bm{x}}_{j}}{\sqrt{d_{v}}}\right)}{\sum_{k=1}^{N}\exp\left(\frac{{\bm{x}}_{i}^{\top}W_{l}^{Q}{W_{l}^{K}}^{\top}{\bm{x}}_{k}}{\sqrt{d_{v}}}\right)}{W_{l}^{V}}^{\top}{\bm{x}}_{j}.$ (19)
We define the expert $f_{j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}^{d_{v}}$, and the score function $s_{i,j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}$ as follows:
$f_{j}(\bm{X})$ $=W^{V^{\top}}_{l}E_{j}\bm{X}=W^{V^{\top}}_{l}\bm{x}_{j},$ (20)
for $i,j=1,\dots,N$. Each output vector ${\bm{h}}_{l,i}\in\mathbb{R}^{d_{v}}$ can then be expressed as an MoE model:
${\bm{h}}_{l,i}$ $=\sum^{N}_{j=1}\frac{\exp(s_{i,j}({\bm{X}}))}{\sum^{N}_{k=1}\exp({s_{i,k}}({\bm{X}}))}\cdot f_{j}({\bm{X}}).$ (22)
This completes the proof of the proposition.
A.2 Proof of Proposition 3.2
Similar to the notation in A.1, we define the new experts and score functions for $j^{\prime}=1,\dots,L$ and $i=1,\dots,N$ as follows:
$f_{N+j^{\prime}}(\bm{X})$ $=W_{l}^{V^{\top}}{\bm{p}}_{j^{\prime}}^{V},$ (23)
$s_{i,N+j^{\prime}}({\bm{X}})$ $=\frac{\bm{X}^{\top}E_{i}^{\top}W_{l}^{Q}W_{l}^{K^{\top}}{\bm{p}}_{j^{\prime}}^{K}}{\sqrt{d_{v}}}=\frac{{\bm{x}}_{i}^{\top}W_{l}^{Q}W_{l}^{K^{\top}}{\bm{p}}_{j^{\prime}}^{K}}{\sqrt{d_{v}}}.$ (24)
Then, using Equation (5), the output of the $l$-th attention head can be expressed as follows:
${\tilde{{\bm{h}}}}_{l}$ $=\mathrm{Attention}\left({\mathbf{X}}^{Q}W_{l}^{Q},\begin{bmatrix}{\mathbf{P}}^{K}\\ {\mathbf{X}}^{K}\end{bmatrix}W_{l}^{K},\begin{bmatrix}{\mathbf{P}}^{V}\\ {\mathbf{X}}^{V}\end{bmatrix}W_{l}^{V}\right)=\left[{\tilde{{\bm{h}}}}_{l,1},\dots,{\tilde{{\bm{h}}}}_{l,N}\right]^{\top}\in\mathbb{R}^{N\times d_{v}},$
${\tilde{{\bm{h}}}}_{l,i}$ $=\sum_{j=1}^{N}\frac{\exp\left(\frac{{\bm{x}}_{i}^{\top}W_{l}^{Q}{W_{l}^{K}}^{\top}{\bm{x}}_{j}}{\sqrt{d_{v}}}\right)}{\mathcal{D}_{i}({\bm{X}})}{W_{l}^{V}}^{\top}{\bm{x}}_{j}+\sum_{j^{\prime}=1}^{L}\frac{\exp\left(\frac{{\bm{x}}_{i}^{\top}W_{l}^{Q}{W_{l}^{K}}^{\top}{\bm{p}}_{j^{\prime}}^{K}}{\sqrt{d_{v}}}\right)}{\mathcal{D}_{i}({\bm{X}})}W_{l}^{V^{\top}}{\bm{p}}_{j^{\prime}}^{V}$
$=\sum_{j=1}^{N}\frac{\exp(s_{i,j}({\bm{X}}))}{\sum_{k=1}^{N+L}\exp(s_{i,k}({\bm{X}}))}f_{j}({\bm{X}})+\sum_{j^{\prime}=1}^{L}\frac{\exp(s_{i,N+j^{\prime}}({\bm{X}}))}{\sum_{k=1}^{N+L}\exp(s_{i,k}({\bm{X}}))}f_{N+j^{\prime}}({\bm{X}}),$ (25)
where $\mathcal{D}_{i}({\bm{X}})=\sum_{j=1}^{N}\exp\left(\frac{{\bm{x}}_{i}^{\top}W_{l}^{Q}{W_{l}^{K}}^{\top}{\bm{x}}_{j}}{\sqrt{d_{v}}}\right)+\sum_{j^{\prime}=1}^{L}\exp\left(\frac{{\bm{x}}_{i}^{\top}W_{l}^{Q}{W_{l}^{K}}^{\top}{\bm{p}}_{j^{\prime}}^{K}}{\sqrt{d_{v}}}\right)$. This completes the proof of the proposition.
Appendix B Comparison to Sparse Mixture of Experts
In this section, we detail how our task-specific prompt pool framework can be viewed as an implementation of a Sparse Mixture of Experts (SMoE) architecture.
By setting the prompt length $L=1$, each prompt in the pool $\mathcal{P}_{t}$ (from Equation (9)) functions as a single prefix expert. We denote these newly introduced experts as $f_{N+1}^{(t)},\dots,f_{N+M}^{(t)}$. While the original $N$ experts from the pre-trained model $f_{1},\dots,f_{N}$ are always active, we apply sparse selection exclusively to these new prefix experts.
Recall from Equation (6) that each head in the MSA layer contains $N$ MoE models ${\tilde{{\bm{h}}}}_{l,1},\dots,{\tilde{{\bm{h}}}}_{l,N}$. A naive selection strategy would apply the $\mathrm{TopK}$ gating function independently to each of these $N$ models. This would require computing $N\times M$ score functions $s_{i,N+j^{\prime}}({\bm{X}})$ for all $i=1,\dots,N$, $j^{\prime}=1,\dots,M$ and could result in a different set of prefix experts being selected for each model.
In contrast, our approach efficiently selects a single, shared set of $K$ new experts across all $N$ MoE models. We achieve this by using an auxiliary score function:
where $i=1,\dots,N$ and $j^{\prime}=1,\dots,M$. Since $\tilde{s}_{i,N+j^{\prime}}({\bm{X}})$ is independent of the index $i$ and depends only on the prompt key $\bm{k}_{j^{\prime}}^{(t)}$, we only need to compute $M$ unique scores. This design significantly reduces the computational burden while enabling the efficient selection of top-$K$ experts from the prompt pool.
While computing the query feature $q({\bm{x}})$ might appear to be an additional expense, this feature is already required by our task predictor (Section 4.3) during both training and inference. Therefore, we reuse this computation for prompt selection at no additional overhead.
Appendix C Comparison of Cascade Voting and the MLP-based Approach
Both the MLP-based head strategy proposed by Le et al. (2025a) and the cascade voting mechanism introduced by Zhou et al. (2024) share the principle of treating each observed relation independently, rather than grouping all relations within a task into a single category. In the MLP-based paradigm, each head corresponds to a specific relation. In contrast, the cascade voting framework scores every relation within a task and selects the one with the highest score as that task’s representative. Subsequently, the representatives from all tasks are assessed, and the task associated with the highest-scoring representative is designated as the predicted task.
A key distinction lies in whether task classification is explicitly trained. The MLP-based paradigm employs a standard supervised approach: an MLP head is allocated to each relation, and the final prediction corresponds to the head with the highest output probability. In contrast, cascade voting dispenses with a trainable classifier, instead leveraging inference-time computations (e.g., distance-based metrics) to identify the most appropriate task.
Regarding memory usage, the MLP-based approach only requires storing relation-specific distributions to facilitate data reconstruction and mitigate catastrophic forgetting. The cascade voting mechanism, however, relies on relation distributions that are further specialized by each prompt pool, resulting in a larger number of stored distributions. Despite this increased memory demand, a notable advantage of cascade voting is that it eliminates the overhead of training a task classifier, thereby expediting the overall training process compared to its MLP-based counterpart.
@@ Appendix C Comparison of Cascade Voting and the MLP-based Approach
Appendix D Impact of the Number of Label Descriptions $D$
To evaluate the sensitivity of WAVE++ to the number of label descriptions $D$, as defined in Equation (11), we conducted experiments on the FewRel and TACRED datasets, varying $D$ from 0 to 7. The results, presented in Table 9, demonstrate that WAVE++ maintains robust performance even with a limited number of descriptions.
These findings indicate that WAVE++ is not heavily dependent on a large set of label descriptions to perform effectively. Indeed, the model achieves high accuracy with only a few descriptions, offering a favorable trade-off between performance and computational efficiency. This robustness makes WAVE++ highly suitable for practical applications where acquiring extensive label descriptions is either infeasible or costly.
Appendix E Gaussian Justification
Figure 6: t-SNE visualization of hidden representations from a representative task in the FewRel dataset.
In WAVE++, we approximate the distribution of prompted representations within each task using a Gaussian model. To empirically validate this approximation, we present a t-SNE visualization of the prompted representations $f_{\theta}({\bm{x}}_{p})$ for a representative task from the FewRel dataset, as shown in Figure 6. The visualization reveals clear clustering patterns in the latent space, suggesting that a Gaussian distribution provides a reasonable approximation of the underlying structure.
To more rigorously assess the validity of the Gaussian assumption, we conduct Royston’s test (Royston, 1992) for multivariate normality. Royston’s test extends the Shapiro–Wilk test to the multivariate setting by aggregating dimension-wise Shapiro–Wilk statistics and comparing the resulting test statistic to a chi-square distribution. The null hypothesis $H_{0}$ assumes that the data are drawn from a multivariate Gaussian distribution, while the alternative hypothesis $H_{1}$ assumes otherwise. We adopt a significance level of $\alpha=0.05$.
Labels 11 15 21 26 53 58 64 75
pp-value 0.87 0.68 0.39 0.09 0.57 0.42 0.65 0.38
Table 10: Royston’s test on samples from eight classes within a task in FewRel, demonstrating that data can be approximated as Gaussian distribution.
The test is performed on a representative FewRel task containing eight relation classes. As reported in Table 10, all classes yield $p$-values greater than $0.05$, indicating no statistically significant evidence against the multivariate normality assumption. These results provide empirical support for the use of Gaussian modeling in our latent space.
More broadly, our objective is to map raw inputs into a representation space where class separation can be effectively achieved using linear decision boundaries. Within this paradigm, Gaussian modeling is a practical and widely adopted choice, as demonstrated in prior continual learning methods such as HiDe-Prompt and EoE. Moreover, we employ a task-specific shared covariance matrix $\bm{\Sigma}^{t}$ across all relations, which substantially reduces the number of parameters and improves scalability without sacrificing empirical performance.
Appendix F Performance under Incorrect Task Prediction